OpenStudy (anonymous):

Determine the length of the arc AB if the diameter of the circle is 10 feet and the angle is 100 degrees. (Use π = 3.14. Round your answer to one decimal place.)

OpenStudy (anonymous):

answer in ft. @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

the formula for arc lenght is 2*pi*r*C/360=2*3.14*(10/2)*100/360=2*3.14*5*100/360=8.722

OpenStudy (alexandervonhumboldt2):

to one decimal place it wold be 8.7

OpenStudy (anonymous):

thanks so much